You observe a 2.0 kg particle moving at a constant speed of 3.6 m/s in a clockwise direction around a circle of radius 4.0 m. (a) What is its angular momentum about the center of the circle? kg·m2/s (b) What is its moment of inertia about an axis through the center of the circle and perpendicular to the plane of the motion? kg·m2 (c) What is the angular velocity of the particle? rad/s
Answer #1

a) Ang. Momentum L = mvr = 2*3.6 *4 =28.8 kg m^2 /s

b) L = Iv/r ==> I = Lr/v = 28.8*4/3.6 = 32 kg m ^2

c) w = v/r = 4/3.6 = 1.111 rad/s